# Grand-Canonical Typicality

### v1

@@ 8 Derivation of Statement 5: Approach to Equilibrium
The first one provides a large part of the first statement in Statement 5: it states that a version of the eigenstate thermalization hypothesis (ETH) suitable for microscopic thermal equilibrium (MITE) implies that the reduced density matrix $\hat{\rho}^{S}_{\Psi_{t}}$ approaches $\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}$ for every initial state $\Psi_{0}\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$. The MITE-ETH was formulated in Statement 5 as (61a) and (61b) and is formulated in Proposition 1 in a more precise version as (108) and (109).
@@
Proposition 1 expresses that for every initial state $\psi_{0}\in\mathbb{S}(\mathscr{H}_{R})$, for most times $t\geq 0$, $\hat{\rho}^{S}_{\psi_{t}}\approx\operatorname{tr}^{B}\hat{\rho}_{R}$. The proof can be found in Appendix D. It is an adaption and generalization of the one given in [43, 34, 11] for the case that $\hat{\rho}_{R}=\hat{\rho}_{\mathrm{mc}}$ and the Hamiltonian has non-degenerate eigenvalues and eigenvalue gaps.
Another proposition concerns the commutation of quantifiers. It is helpful to introduce the symbol $\bigvee\hskip-8.64961pt|\,\,$} to denote “for most.” If $A(x,y)$ is any statement about objects $x$ and $y$, then $\forall x\forall y:A(x,y)$” is equivalent to $\forall y\forall x:A(x,y)$.” Likewise, ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x{\textstyle\bigvee\hskip-8.64961pt|\,\,}y:A(x,y)$” is equivalent to ${\textstyle\bigvee\hskip-8.64961pt|\,\,}y{\textstyle\bigvee\hskip-8.64961pt|\,\,}x:A(x,y)$[13, Footnote 7] if both ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}y$ refer to probability spaces (i.e., normalized measures); the situation is more subtle when talking about most times $t\in[0,\infty)$ because the uniform measure on $0,\infty)$ is not normalizable. In fact, ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x{\textstyle\bigvee\hskip-8.64961pt|\,\,}t:A(x,t)$” implies that ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t{\textstyle\bigvee\hskip-8.64961pt|\,\,}x:A(x,t)$” but in general not vice versa [50, Footnote 5]. The following proposition provides a sufficient condition for commuting ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ from the left to the right of ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x$.
@@
The proof can be found in Appendix E.
@@
The proof can be found in Appendix F.
Derivation of Statement 5. We assume that $\hat{H}$ satisfies the ETH (108)–(109), and that $D_{G}$ is not too large. Then Proposition 1 for $\mathscr{H}_{R}=\mathscr{H}_{\mathrm{gmc}}$ yields that for every $\Psi_{0}\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$, for most times $t\in[0,\infty)$, $\hat{\rho}_{\Psi_{t}}^{S}\approx\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}$. Statement 1a tells us that $\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}_{\mathrm{gG}}^{S}$. We thus obtain the first statement (62) of Statement 5.
We now turn to the conditional wave function. Theorem 2 in [14] shows that for every $t$, most orthonormal bases $b$ of $\mathscr{H}^{S^{c}}$ are such that $\mathscr{L}_{\psi^{S}(t)}\approx\mathrm{GAP}_{\hat{\rho}^{S}_{\Psi_{t}}}$. By the continuous dependence (91) of $\mathrm{GAP}_{\hat{\rho}}$ on $\hat{\rho}$, $\mathrm{GAP}_{\hat{\rho}^{S}_{\Psi_{t}}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$ for most $t$ and regardless of the basis $b$. Thus,
@@
We want to move the ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$” to the left of ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$” because we want to apply the same basis $b$ at all times $t$. As discussed before Proposition 2, even commuting ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$ requires further assumptions, to which we turn now. We express that $\mathscr{L}_{\psi^{S}(t)}$ is close to $\mathrm{GAP}_{\hat{\rho}^{S}_{\mathrm{gG}}}$ by means of integrating them against arbitrary “test” functions $f:\mathbb{S}(\mathscr{H}^{S})\to\mathbb{R}$ as in (93). Proposition 2 for $X=ONB(\mathscr{H}^{S^{c}})$, $x=b$, and $F(b,t)=|\mathscr{L}_{\psi^{S}(t)}(f)-\mathrm{GAP}_{\hat{\rho}^{S}_{\mathrm{gG}}}(f)|$ (which satisfies $0\leq F(b,t)\leq C=2\|f\|_{\infty}$) allows us to interchange ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$, as its hypothesis (112) is satisfied by virtue of Proposition 3 with $g(\Psi)=|\mathscr{L}_{\psi^{S}}(f)-\mathrm{GAP}_{\hat{\rho}^{S}_{\mathrm{gG}}}(f)|$ (note that $\psi^{S}$ depends on $\Psi$) for arbitrary but fixed $b$ and $f$; after all, $F(b,t)=g(\Psi_{t})$. Thus,
@@
An alternative strategy of the derivation starts from the fact, mentioned above, that $\forall\Psi_{0}\,\forall t\geq 0\,{\textstyle\bigvee\hskip-8.64961pt|\,\,}b:\>\mathscr{L}_{\psi^{S}(t)}\approx\mathrm{GAP}_{\hat{\rho}_{\Psi_{t}}^{S}}$. Now we try first, before replacing $\hat{\rho}_{\Psi_{t}}^{S}$ by $\hat{\rho}_{\mathrm{gG}}^{S}$, to move ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$” to the left of the $t$-quantifier. It must still be expected that $\forall t$” has to be replaced by the weaker ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$” because it may well happen that every $b$ has some bad, exceptional $t$’s [13]. We apply Proposition 3 to $g(\Psi)=|\mathscr{L}_{\psi^{S}}(f)-\mathrm{GAP}_{\hat{\rho}_{\Psi}^{S}}(f)|$, then Proposition 2 to $X=ONB(\mathscr{H}^{S^{c}})$, $x=b$, and $F(b,t)=|\mathscr{L}_{\psi^{S}(t)}(f)-\mathrm{GAP}_{\hat{\rho}_{\Psi_{t}}^{S}}(f)|=g(\Psi_{t})$, so we can interchange ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$. Now if each of two statements $s_{1}(t)$ and $s_{2}(t)$ is true for $(1-\varepsilon)$-most $t$, then $s_{1}(t)$ and $s_{2}(t)$” is true for at least $(1-2\varepsilon)$-most $t$. Since ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t:\hat{\rho}_{\Psi_{t}}^{S}\approx\hat{\rho}_{\mathrm{gG}}^{S}$ by (62) and for those $t$, $\mathrm{GAP}_{\hat{\rho}_{\Psi_{t}}^{S}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$ by (91), we obtain (115), as claimed.
@@ Appendix D Proof of Proposition 1
Appendix E Proof of Proposition 2
@@
Appendix F Proof of Proposition 3
@@
Acknowledgments. C.V. was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – TRR 352 – Project-ID 470903074. Moreover, C.V. acknowledges financial support by the ERC Starting Grant “FermiMath” No. 101040991 and the ERC Consolidator Grant “RAMBAS” No. 10104424, funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 8 Derivation of Statement 5: Approach to Equilibrium
The first one provides a large part of the first statement in Statement 5: it states that a version of the eigenstate thermalization hypothesis (ETH) suitable for microscopic thermal equilibrium (MITE) implies that the reduced density matrix $\hat{\rho}^{S}_{\Psi_{t}}$ approaches $\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}$ for every initial state $\Psi_{0}\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$. The MITE-ETH was formulated in Statement 5 as (65a) and (65b) and is formulated in Proposition 2 in a more precise version as (123) and (124).
@@
Proposition 2 expresses that for every initial state $\psi_{0}\in\mathbb{S}(\mathscr{H}_{R})$, for most times $t\geq 0$, $\hat{\rho}^{S}_{\psi_{t}}\approx\operatorname{tr}^{B}\hat{\rho}_{R}$. The proof can be found in Appendix E. It is an adaption and generalization of the one given in [48, 39, 14] for the case that $\hat{\rho}_{R}=\hat{\rho}_{\mathrm{mc}}$ and the Hamiltonian has non-degenerate eigenvalues and eigenvalue gaps.
Another proposition concerns the commutation of quantifiers. It is helpful to introduce the symbol $\bigvee\hskip-8.64961pt|\,\,$} to denote “for most.” If $A(x,y)$ is any statement about objects $x$ and $y$, then $\forall x\forall y:A(x,y)$” is equivalent to $\forall y\forall x:A(x,y)$.” Likewise, ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x{\textstyle\bigvee\hskip-8.64961pt|\,\,}y:A(x,y)$” is equivalent to ${\textstyle\bigvee\hskip-8.64961pt|\,\,}y{\textstyle\bigvee\hskip-8.64961pt|\,\,}x:A(x,y)$[16, Footnote 7] if both ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}y$ refer to probability spaces (i.e., normalized measures); the situation is more subtle when talking about most times $t\in[0,\infty)$ because the uniform measure on $0,\infty)$ is not normalizable. In fact, ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x{\textstyle\bigvee\hskip-8.64961pt|\,\,}t:A(x,t)$” implies that ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t{\textstyle\bigvee\hskip-8.64961pt|\,\,}x:A(x,t)$” but in general not vice versa [56, Footnote 5]. The following proposition provides a sufficient condition for commuting ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ from the left to the right of ${\textstyle\bigvee\hskip-8.64961pt|\,\,}x$.
@@
The proof can be found in Appendix F.
@@
The proof can be found in Appendix G.
Derivation of Statement 5. We assume that $\hat{H}$ satisfies the ETH (123)–(124), and that $D_{G}$ is not too large. Then Proposition 2 for $\mathscr{H}_{R}=\mathscr{H}_{\mathrm{gmc}}$ yields that for every $\Psi_{0}\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$, for most times $t\in[0,\infty)$, $\hat{\rho}_{\Psi_{t}}^{S}\approx\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}$. Statement 1a tells us that $\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}_{\mathrm{gG}}^{S}$. We thus obtain the first statement (66) of Statement 5.
We now turn to the conditional wave function. Theorem 2 in [17] shows that for every $t$, most orthonormal bases $b$ of $\mathscr{H}^{\overline{S}}$ are such that $\mathscr{L}_{\psi^{S}(t)}\approx\mathrm{GAP}_{\hat{\rho}^{S}_{\Psi_{t}}}$. By the continuous dependence (99) of $\mathrm{GAP}_{\hat{\rho}}$ on $\hat{\rho}$, $\mathrm{GAP}_{\hat{\rho}^{S}_{\Psi_{t}}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$ for most $t$ and regardless of the basis $b$. Thus,
@@
We want to move the ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$” to the left of ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$” because we want to apply the same basis $b$ at all times $t$. As discussed before Proposition 3, even commuting ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$ requires further assumptions, to which we turn now. We express that $\mathscr{L}_{\psi^{S}(t)}$ is close to $\mathrm{GAP}_{\hat{\rho}^{S}_{\mathrm{gG}}}$ by means of integrating them against arbitrary “test” functions $f:\mathbb{S}(\mathscr{H}^{S})\to\mathbb{R}$ as in (101). Proposition 3 for $X=ONB(\mathscr{H}^{\overline{S}})$, $x=b$, and $F(b,t)=|\mathscr{L}_{\psi^{S}(t)}(f)-\mathrm{GAP}_{\hat{\rho}^{S}_{\mathrm{gG}}}(f)|$ (which satisfies $0\leq F(b,t)\leq C=2\|f\|_{\infty}$) allows us to interchange ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$, as its hypothesis (127) is satisfied by virtue of Proposition 4 with $g(\Psi)=|\mathscr{L}_{\psi^{S}}(f)-\mathrm{GAP}_{\hat{\rho}^{S}_{\mathrm{gG}}}(f)|$ (note that $\psi^{S}$ depends on $\Psi$) for arbitrary but fixed $b$ and $f$; after all, $F(b,t)=g(\Psi_{t})$. Thus,
@@
An alternative strategy of the derivation starts from the fact, mentioned above, that $\forall\Psi_{0}\,\forall t\geq 0\,{\textstyle\bigvee\hskip-8.64961pt|\,\,}b:\>\mathscr{L}_{\psi^{S}(t)}\approx\mathrm{GAP}_{\hat{\rho}_{\Psi_{t}}^{S}}$. Now we try first, before replacing $\hat{\rho}_{\Psi_{t}}^{S}$ by $\hat{\rho}_{\mathrm{gG}}^{S}$, to move ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$” to the left of the $t$-quantifier. It must still be expected that $\forall t$” has to be replaced by the weaker ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$” because it may well happen that every $b$ has some bad, exceptional $t$’s [16]. We apply Proposition 4 to $g(\Psi)=|\mathscr{L}_{\psi^{S}}(f)-\mathrm{GAP}_{\hat{\rho}_{\Psi}^{S}}(f)|$, then Proposition 3 to $X=ONB(\mathscr{H}^{\overline{S}})$, $x=b$, and $F(b,t)=|\mathscr{L}_{\psi^{S}(t)}(f)-\mathrm{GAP}_{\hat{\rho}_{\Psi_{t}}^{S}}(f)|=g(\Psi_{t})$, so we can interchange ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t$ and ${\textstyle\bigvee\hskip-8.64961pt|\,\,}b$. Now if each of two statements $s_{1}(t)$ and $s_{2}(t)$ is true for $(1-\varepsilon)$-most $t$, then $s_{1}(t)$ and $s_{2}(t)$” is true for at least $(1-2\varepsilon)$-most $t$. Since ${\textstyle\bigvee\hskip-8.64961pt|\,\,}t:\hat{\rho}_{\Psi_{t}}^{S}\approx\hat{\rho}_{\mathrm{gG}}^{S}$ by (66) and for those $t$, $\mathrm{GAP}_{\hat{\rho}_{\Psi_{t}}^{S}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$ by (99), we obtain (130), as claimed.
@@ Appendix D Proof of Proposition 1
Existence: By (85),
$\hat{I}^{a}\otimes\hat{Q}^{bc}-\hat{I}^{a}\otimes\hat{I}^{b}\otimes\hat{Q}^{c}=\hat{Q}^{ab}\otimes\hat{I}^{c}-\hat{Q}^{a}\otimes\hat{I}^{b}\otimes\hat{I}^{c}\,.$ (143)
Since the left-hand side is of the form $\hat{I}^{a}\otimes\hat{R}^{bc}$ and the right-hand side of the form $\hat{R}^{ab}\otimes\hat{I}^{c}$, both sides have to be of the form $\hat{I}^{a}\otimes\hat{Q}^{b}\otimes\hat{I}^{c}$. For the left-hand side, it follows (after dropping a factor $\hat{I}^{a}$) that $\hat{Q}^{bc}=\hat{Q}^{b}\otimes\hat{I}^{c}+\hat{I}^{b}\otimes\hat{Q}^{c}$. For the right-hand side, it follows (after dropping a factor $\hat{I}^{c}$) that $\hat{Q}^{ab}=\hat{Q}^{a}\otimes\hat{I}^{b}+\hat{I}^{a}\otimes\hat{Q}^{b}$, as claimed.
Uniqueness: Insert (86a) into (85), so $\hat{I}^{a}\otimes\hat{I}^{b}\otimes\hat{Q}^{c}$ cancels out; drop the common factor $\hat{I}^{c}$ and take the partial trace over $\mathscr{H}^{a}$ to obtain that $\hat{Q}^{b}=(\dim\mathscr{H}^{a})^{-1}(\operatorname{tr}^{a}\hat{Q}^{ab}-(\operatorname{tr}\hat{Q}^{a})\hat{I}^{b})$.
Appendix E Proof of Proposition 2
@@
Appendix F Proof of Proposition 3
@@
Appendix G Proof of Proposition 4
@@
Acknowledgments. We thank an audience of the Rutgers University mathematical physics webinar for helpful discussion. C.V. was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – TRR 352 – Project-ID 470903074. Moreover, C.V. acknowledges financial support by the ERC Starting Grant “FermiMath” No. 101040991 and the ERC Consolidator Grant “RAMBAS” No. 10104424, funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.